Suppose f(x,y)=1x^2+3xy−2y^2, P=(3,2), and u=(−5/13,12/13).

A. Compute the gradient of f.
∇f=___i+___j
Note: Your answers should be expressions of x and y; e.g. "3x - 4y"

B. Evaluate the gradient at the point P.
(∇f)(3,2)=___i+___j
Note: Your answers should be numbers

C. Compute the directional derivative of f at P in the direction u .
(Duf)(P)=

Note: Your answer should be a number
1
Expert's answer
2020-10-19T17:42:50-0400

A. The gradient of "f(x,y)" is given by "∇f(x,y)=f_x(x,y) \vec{i}+f_y(x,y) \vec{j}."

"f(x,y)=1x^2+3xy\u22122y^2,"

"f_x(x,y)=2x+3y,"

"f_y(x,y)=3x-4y."

So we have:

"∇f(x,y)=(2x+3y) \vec{i}+(3x-4y) \vec{j}."


Answer. "∇f=(2x+3y) \vec{i}+(3x-4y) \vec{j}."


B. "∇f(x_0,y_0)=f_x(x_0,y_0) \vec{i}+f_y(x_0,y_0) \vec{j}." Hence at the point "P(3,2)" we have:

"∇f(3,2)=(2 \cdot 3+3 \cdot 2) \vec{i}+ (3 \cdot 3-4 \cdot 2) \vec{j}= 12 \vec{i} +1 \vec{j}."


Answer. "∇f(3,2)= 12 \vec{i} +1 \vec{j}."


C. "D_u f(P)= ∇f(3,2) \cdot \vec{u}," where "\vec{u}" is a unit vector. Let's check it.

"||u||=\sqrt{(-\frac{5}{13})^2+(\frac{12}{13})^2}=\sqrt{\frac{25}{169}+\frac{144}{169}}=1."

So "\vec{u}=-\frac{5}{13} \vec{i}+\frac{12}{13} \vec{j}" is a unit vector.

"D_u f(3, 2)=(12 \vec{i} +1 \vec{j}) \cdot (-\frac{5}{13} \vec{i}+\frac{12}{13} \vec{j})=
12 \cdot (-\frac{5}{13})+1 \cdot (\frac{12}{13})="

"=-\frac{60}{13}+ \frac{12}{13}= - \frac{48}{13}=-3 \frac{9}{13}."


Answer. "D_u f(P) = -3 \frac{9}{13}."